PS1 IPP Czar Logs for the week 2011.03.21 - 2011.03.27

(Up to PS1 IPP Czar Logs)

Monday : 2011.03.21

Tuesday : 2011.03.22

Wednesday : 2011.03.23

Roy is czar.

  • 08:00 No data last night....
  • 13:15 Bill is doing some processing from a pantasks on ippc25.

Thursday : 2011.03.24

Roy is czar.

  • 08:00 No data last night....
  • 14:00 CZW Stopping processing for summit_id addition to database.
  • 14:41 CZW Cleanup finally finished, beginning dumps of GPC1 and ISP databases.
  • 15:28 CZW Waiting for initial GPC1 dump to finish. Rebuilding tag for ippuser.
  • 15:33 CZW Dump finished. Making changes to GPC1 database for schema_version 1.1.69.
  • 15:45 CZW New IPP user build complete.
  • 18:13 CZW Commit transaction, completing upgrade to schema 1.1.69. Beginning second dump.
  • 19:35 CZW Second dump finished. Starting on ISP (which should be faster?).
  • 19:55 CZW ISP upgrade aborted. I can't get the read lock, and I've been waiting almost a half hour. I have removed the add.database isp entries from summitcopy/input and registration/input.
  • 19:57 CZW Restarting servers: ./ start.server / start / run
  • 20:06 CZW bugfix on pztool due to missing command line options.
  • 20:19 CZW commented out offending line in pztool.c (511) because I can't see why it's still complaining.
  • 20:28 CZW fixing new bug at the pztool -advance stage. Luckily, this should be the end of where changes happened.

Friday : 2011.03.25

  • 05:00 EAM ppStack was failing -- error in the new timers; I updated the code and merged the changes back to the trunk
  • 06:00 EAM ppSub is failing for MD06 y-band stack-stack diffs. I've set the run (diff_id = 117893) state to wait until it can be debugged
  • 09:50 EAM 1 diff failed due to a corrupt warp. I reverted with --redirect-output --warp_id 175863 --skycell_id skycell.1805.128
  • 09:55 EAM 6 warps failed from I/O, I reverted manually; 2 other diffs had failed due to a negative kron normalization (implies a very bad subtraction) : I merged 2 lines of code that address this as bad quality.
  • 10:30 EAM needed to re-run one camRun due to corrupted masks crashing warp: perl ~/src/ipp-20110218/tools/ --cam_id 187761

Saturday : 2011.03.26

Sunday : 2011.03.27